# Detection of disk-jet co-precession in a tidal disruption event

### v1

@@
AT2020afhd (a.k.a. ZTF20abwtifz) is an optical transient situated at the nucleus of the galaxy LEDA 145386 (Fig. S1), at a redshift of 0.027 (?, ?). The transient was discovered by the Zwicky Transient Facility (ZTF) in 2020, with a single detection at a $g$-band magnitude of $\sim 20$. On January 4, 2024, ZTF detected a significant re-brightening (?), with a peak magnitude of $g_{\rm ABmag}\sim 16.6$. Follow-up optical spectra reveal the presence of a blue continuum (consistent with the optical colors of $g-r\sim-0.02$), an He ii emission line and broad Balmer emission, leading to the classification of the re-brightening as a tidal disruption event (TDE, (?)). Moreover, the optical decline rate follows approximately a $t^{-5/3}$ power law (see Fig. 1), consistent with the debris fallback rate predicted by TDE theories (?, ?, ?). Additionally, we determined the central black hole (BH) mass using two independent methods (Supplementary 7), both of which yielded consistent results. For the subsequent analysis in this work, we adopted $\log(M_{\rm BH}/M_{\odot})=6.7\pm 0.5$.
@@
In X-rays, several Neil Gehrels Swift Observatory (Swift) monitoring programs commenced since January 26, 2024 (25 days after the re-brightening). These observations revealed significant variations on timescales of approximately 25–40 days, during which the peak luminosities exceeded the dips by more than one order of magnitude (see Fig. 1). Such variations are also evident in our Neutron star Interior Composition ExploreR (NICER) campaign, triggered 10 days after the initial Swift program. The peak X-ray luminosity is over two orders of magnitude higher that in 2020, suggesting the variations are associated with the newly occurred TDE. Moreover, during the first 300 days, the X-ray spectrum is ultra soft and can be well described by a multi-color disk blackbody component with $kT_{\rm in}\sim 10^{5.7-6.4}$~K. The derived blackbody temperature closely follows the evolution of the X-ray luminosity, increasing with rising luminosity and decreasing as it declines (see Fig. 2a).
@@
By 215 days after the re-brightening, the 0.3–2 keV X-ray luminosity had dropped by over an order of magnitude and began exhibiting clear periodic variations, visible to the naked eye. The variation amplitude remained above an order of magnitude. A Lomb–Scargle periodogram (LSP) of background-subtracted X-ray data from 215–294 days (August 3 to October 21) identifies a period of $19.6\pm 1.5$ days with a statistical significance exceeding $3.79\sigma$ (see Supplementary for details). Notably, the short-term modulation is absent in the high-cadence optical and UV photometry. The consistent periodic behavior observed in both the luminosity and blackbody temperature is reminiscent of the TDE AT2020ocn (?). In a similar case, AT2020ocn exhibited X-ray modulations with a 15-day period lasting approximately 130 days, which were attributed to Lense-Thirring (LT) precession, although the possibility of radiation-pressure instabilities could not be entirely excluded.
@@
Compared to other TDEs with early radio detections (within a hundred days of discovery), AT2020afhd exhibited unprecedented high-amplitude variations in radio bands on short-term timescales (as shown in Fig. 3), pointing to the emergence of a new class of radio TDEs. Our VLBA U- and C-band observations across different epochs confirmed that the emitting region remained point-like (Fig. S2). The consistent flux densities observed with VLBA, VLA, and e-MERLIN at similar times indicate that the radio emission, at least at 5–6 GHz, originated from a compact, unresolved source, effectively ruling out angular resolution effects. Additionally, flux calibration offsets among the VLA, e-MERLIN and ATCA observations, expected to be up to 22% (Supplementary 1.3.5), cannot account for the observed variations. The effect of interstellar scintillation (?), which would produce hour-scale variations, is excluded based on the consistent flux densities detected in multiple VLA observations taken within one hour (Supplementary 8.2). Emission from winds or jet-wind interactions with the circumnuclear medium is expected to be isotropic and should not exhibit significantly modulated variations. Therefore, the previously uncharted short-term radio variations observed in AT2020afhd are most likely driven by the dynamic evolution of jets.
@@
Unfortunately, the sparse sampling of radio observations, combined with contamination from non-periodic variations on both long timescales (hundreds of days) and short timescales (days), makes it challenging to directly extract periodicity from the radio data and reliably assess its significance through comprehensive analysis. Instead, we evaluated the connection between X-ray and radio variations using the discrete cross-correlation function (?) (Supplementary 3). This analysis revealed a significant $4.26\sigma$ correlation between the X-ray and radio emissions, with a time lag of $19.0^{+0.7}_{-0.6}$, consistent with the period of the X-ray variations (see Figs. 4b and c). The coincidence between the negative time lag and the period suggests that the jet may have stably formed earlier than the disk. The nearly synchronized radio and X-ray variations suggest a common mechanism tightly regulating both emissions. The tens-of-day period aligns with disk precession in TDEs (?, ?), making synchronized disk-jet precession the most natural explanation. While supported by general relativistic magnetohydrodynamics (GRMHD) simulations (?, ?), this phenomenon remains observationally unconfirmed. The only other known case of simultaneous X-ray and radio quasi-periodic oscillations (QPOs) is the microquasar GRS 1915+105, which occurs on timescales of 20–50 minutes (?, ?) and has been attributed to disk radiation-pressure instabilities (e.g., (?, ?)).
@@
To test the disk-jet precession scenario, we first introduced a rigid-body LT precession model under conditions where the accretion rate remained above the Eddington limit during the first year of a TDE involving a SMBH with a mass of $10^{6.7\pm 0.5}~M_{\odot}$ disrupted a solar-like star. During this period, the disk remains geometrically thick, characterized by a disk angular semi-thickness $H/R>\alpha$, where $\alpha$ is the disk viscosity parameter (?), $H$ is the disk thickness, and $R$ is the disk radius. The disk is assumed to extend from the ISCO to the debris circularization radius (?, ?), and the amplitude of the X-ray variations is attributed to changes in the disk’s projected area and periodic obscuration by the outer disk (see Supplementary 8.1 for details). The BH spin parameter can then be estimated by assuming that the observed period $T_{\rm prec}\sim 19.6$ days corresponds to the LT precession period. Adopting several power-law indices for the surface density profile (?), and a BH mass range of $M_{\bullet}=10^{6.7\pm 0.5}~M_{\odot}$, we found that the spin parameter $a_{\bullet}$ fell within the range of (0.460.14)-(0.46-0.14) or 0.110.350.11-0.35 (Fig. S8). Since a negative spin parameter would result in a larger ISCO and, consequently, a larger inner disk radius, it would be challenging to explain the observed hot disk emission and could even hinder the occurrence of the TDE. Therefore, a positive spin parameter was favored. Eventually, with this model we are able to reasonably reproduce the observed variations and constrain the observer’s viewing angle relative to the BH spin to $\theta_{\rm obs}\sim 38.4^{+0.5\circ}_{-0.6}$, and the disk precession angle to $\theta_{\rm i}\sim 14.5^{\circ}\pm 0.5^{\circ}$ (see the schematics in Fig. 5a).
@@
We then examined the jet precession scenario, in which the radio luminosity varies as the jet cone shifts toward and away from the line of sight. The peak-to-dip ratio is determined by the Doppler factor ratio, which depends on the jet Lorentz factor ($\Gamma$) and the angle between the observer and the jet axis (Supplementary). The luminosity peaks when the difference between $\theta_{\rm obs}$ and $\theta_{\rm i}$ is minimized and dips when it is maximized (see Fig. 5a). Adopting $\theta_{\rm obs}\sim 38.4^{\circ}$ and $\theta_{\rm i}\sim 14.5^{\circ}$, we derived $\Gamma\leq 1.6$. Jets can be powered by two primary mechanisms: the Blandford-Znajek (BZ; (?)) mechanism, driven by BH spin, and the Blandford-Payne (BP; (?)) mechanism, powered by disk rotation. Both mechanisms can account for our radio observations, with the BZ mechanism requiring a stronger magnetic field than the BP mechanism, i.e., $B_{\bullet}\sim 2.8\times 10^{3}\rm~G$ for the former and $B_{\bullet}\sim 1.5\times 10^{3}\rm~G$ for the latter. In Fig. 5b, we present a comparison between our disk-jet precession model and the observations.
@@
Figure 2: Temporal evolution of spectral parameters. a, The diskbb temperature derived from X-ray spectra. The luminosity evolution is depicted with grey symbols to illustrate its correlation with temperature; the two parameters generally co-evolve, with the temperature reaching its peak alongside luminosity during the first 300 days. b, The in-band spectral index ($F\propto\nu^{\alpha}$) derived from the VLA (4–8 GHz) and ATCA (4–10 GHz) data. The stars indicate the VLA observations used for the radio SED modeling (Supplementary). c, The peak frequency, $\nu_{\rm p}$, and the peak flux, $F_{\rm p}$, derived from the SED modeling.
Figure 3: Comparison of TDEs with early intense radio detection. The luminosities of the two on-axis jetted TDEs, Swift J1644+57 and AT2022cmc, have been rescaled for clarity. The AMI-LA data at 15.5 GHz for Swift J1644+57, ASSASN–14li, AT2019azh and AT2022cmc were adapted from (?), (?), (?), and (?), respectively.
Figure 4: Timing analysis of the X-ray and radio data. a, Lomb–Scargle periodogram of the X-ray lightcurve. The calculation included data taken between August 3 and October 21, during which the X-ray quasi-periodic variations were clearly present (see Supplementary). b, Cross-correlation function between the X-ray and radio data. The data used in this calculation are indicated by the gray-shaded region in Figs. 1a and b. The histogram represents the distribution obtained from bootstrap simulations, with the red dashed line marking the median of the distribution at 19.0-19.0 days. c, Folded X-ray and radio lightcurves with a period of 19.6 days. The radio data were rebinned into 0.1-phase intervals using a weighted mean for clarity. The data used in this calculation are marked in Fig. 5b.
Figure 5: The disk-jet precession model. a, Schematics of the proposed disk-jet precession model. $\theta_{\rm obs}$ and $\theta_{\rm i}$ represent the viewing angle of the system and the disk/jet precession angle around the black hole axis, respectively. The left and right plots correspond to the phases of the X-ray and radio variations when the luminosity is relatively low and high, respectively. b, Comparison of the disk-jet precession model (the lower and upper black curves) with X-ray (0.1–2 keV) and radio (5–6 GHz) observations. In the presented model, we adopted a BH mass of $M_{\bullet}=10^{6.7}M_{\odot}$, a scaleheight ratio of $H/R\sim 1$, and an outer disk radius equal to the circularization radius of the debris. As a result, we determined an inclination angle of $\theta_{\rm obs}\sim 37.8-38.9^{\circ}$, a disk/jet precession angle of $\theta_{\rm i}\sim 14-15^{\circ}$, and a Doppler factor of $\Gamma\leq 1.6$.

### v2

@@
Detection of disk-jet co-precession
A disrupted star reveals a wobbling black hole disk and jet, seen in rhythmic X-ray and radio pulses every 19.6 days.
@@
Introduction
Tidal disruption events (TDEs) are transient phenomena that occur when a star ventures too close to a supermassive black hole (SMBH) and is torn apart by its tidal forces (e.g., (?, ?)). The resulting stellar debris falls back toward the SMBH, forming a nascent accretion disk and, in some cases, launching (mildly) relativistic jets (e.g., (?, ?, ?, ?, ?)). This process unfolds over timescales of months to years, offering a unique opportunity to study accretion and jet-launching physics around SMBHs in real time.
In TDEs, the angular momentum of the accretion disk, imparted by the disrupted star, is often misaligned with the spin axis of the central Kerr black hole (BH). This misalignment is expected to induce Lense-Thirring (LT, (?)) precession of the disk and associated jet, driven by frame-dragging effects in the strong-field regime of general relativity (e.g., (?, ?)). General relativistic magnetohydrodynamic (GRMHD) simulations support this picture, predicting coupled disk-jet precession (e.g., (?, ?)); however, direct observational confirmation remains elusive. Thus far, disk or jet precession has only been observed separately in various accreting systems, as evidenced by either X-ray (e.g., (?, ?, ?)) or radio observations (e.g., (?, ?)).
The detection of disk-jet co-precession has been impeded by several observational challenges, including the transient nature of the disk and jet, limited cadence in radio monitoring, viewing angle effects, and contamination from other sources of variability. Here, we present the most compelling evidence to date for disk-jet co-precession in a recent TDE, enabled by exceptionally dense temporal coverage in both X-ray and radio observations.
Results
Optical counterpart
@@ Results
AT2020afhd (a.k.a. ZTF20abwtifz) is an optical transient situated at the nucleus of the galaxy LEDA 145386 (fig. S1), at a redshift of 0.027 (?, ?). The transient was discovered by the Zwicky Transient Facility (ZTF) in 2020, with initial detections at a $g$-band magnitude of $\sim 20$. On January 4, 2024, ZTF detected a significant re-brightening (?), with a peak magnitude of $g_{\rm ABmag}\sim 16.6$. Follow-up optical spectra reveal the presence of a blue continuum (consistent with the optical colors of $g-r\sim-0.02$), an He ii emission line and broad Balmer emission, leading to the classification of the re-brightening as a TDE (?). Moreover, the optical decline rate follows approximately a $t^{-5/3}$ power-law (see Fig. 1), consistent with the debris fallback rate predicted by TDE theories (?, ?, ?). Additionally, we determined the central BH mass using two independent methods (Supplementary 5), both of which yielded consistent results. For the subsequent analysis in this work, we adopt a BH mass of $\log(M_{\rm BH}/M_{\odot})=6.7\pm 0.5$ and take MJD 60310 as the initial date of the re-brightening.
X-ray counterpart
@@
In X-rays, several Neil Gehrels Swift Observatory (Swift) monitoring programs commenced since January 26, 2024 (25 days after the re-brightening). These observations revealed significant variations on timescales of approximately 25–40 days, during which the peak luminosities exceeded the dips by more than one order of magnitude (see Fig. 1). Such variations are also evident in our Neutron star Interior Composition ExploreR (NICER) campaign, triggered 10 days after the initial Swift program. The peak X-ray luminosity is approximately two orders of magnitude higher than in 2020, suggesting the variations are associated with the newly occurred TDE. Moreover, during the first 300 days, the X-ray spectrum remains ultrasoft, dominated by a multi-color disk blackbody component with $kT_{\rm in}\sim 10^{5.7-6.4}$~K. The derived blackbody temperature closely follows the evolution of the X-ray luminosity, increasing with rising luminosity and decreasing as it declines (see Fig. 2A).
@@
By 215 days after the re-brightening, the 0.3–2 keV X-ray luminosity had dropped by over an order of magnitude and began exhibiting clear periodic variations, visible to the naked eye. The variation amplitude remained above an order of magnitude. A Lomb-Scargle periodogram (LSP) of background-subtracted X-ray data from 215–294 days (August 3 to October 21) identifies a period of $19.6\pm 1.5$ days with a statistical significance exceeding $3.79\sigma$ (Fig. 3A, see Materials and Methods for details). Notably, the short-term modulation is absent in the high-cadence optical and UV photometry. The consistent periodic behavior observed in both the luminosity and blackbody temperature is reminiscent of the TDE AT2020ocn (?). In a similar case, AT2020ocn exhibited X-ray modulations with a 15-day period lasting approximately 130 days, which were attributed to LT precession, although the possibility of radiation-pressure instabilities could not be entirely excluded.
Radio counterpart
@@
Compared to other TDEs with early radio detections (within a hundred days of discovery), AT2020afhd exhibited unprecedented high-amplitude variations in radio bands on short-term timescales (as shown in Fig. 4), pointing to the emergence of a new class of radio TDEs. Our VLBA U- and C-band observations across different epochs confirmed that the emitting region remained point-like (fig. S2). The consistent flux densities observed with VLBA, VLA, and e-MERLIN at similar times indicate that the radio emission, at least at 5–6 GHz, originated from a compact, unresolved source, effectively ruling out angular resolution effects. Flux calibration offsets among the VLA, e-MERLIN and ATCA observations, expected to be up to 22% (Supplementary 1.3.5), cannot account for the observed variations. The effect of interstellar scintillation (?), which would produce hour-scale variations, is excluded based on the consistent flux densities detected in multiple VLA observations taken within one hour (Fig. 1B). Emission resulting from winds or jet-wind interactions with the circumnuclear medium is expected to be isotropic and should not exhibit significant modulation on timescales of tens of days. Therefore, the previously uncharted short-term radio variations observed in AT2020afhd are most likely driven by the dynamic evolution of jets.
Cross-correlation between X-ray and radio emission
@@
Unfortunately, the sparse sampling of radio observations, combined with contamination from non-periodic variations on both long timescales (hundreds of days) and short timescales (days), makes it challenging to directly extract periodicity from the radio data and reliably assess its significance through comprehensive analysis. Instead, we evaluated the connection between X-ray and radio variations using the discrete cross-correlation function ( (?), see Materials and Methods). This analysis revealed a significant $4.26\sigma$ correlation between the X-ray and radio emissions, with a primary peak at a time lag of $19.0^{+0.7}_{-0.6}$ days (Figs. 3B and C). In addition, two secondary peaks are found at lags of 0 and 40-40 days, corresponding to integer multiples of the X-ray variability period. These findings suggest that the primary lag of 19-19 days may result from the limited temporal coverage of the data and noise across different timescales, rather than indicating a definitive physical lead of the radio emission. The nearly synchronized radio and X-ray variations suggest a common mechanism tightly regulating both emissions. The tens-of-day period aligns with disk precession in TDEs (?, ?), making synchronized disk-jet precession the most natural explanation.
Disk-jet co-precession model
@@
To test the disk-jet precession scenario, we first introduced a rigid-body LT precession model under conditions where the accretion rate remained above the Eddington limit during the first year of a TDE involving a SMBH with a mass of $10^{6.7\pm 0.5}~M_{\odot}$ disrupted a solar-like star. During this period, the disk remains geometrically thick, characterized by a disk angular semi-thickness $H/R>\alpha$, where $\alpha$ is the disk viscosity parameter (?), $H$ is the disk thickness, and $R$ is the disk radius. The disk is assumed to extend from the ISCO to the debris circularization radius (?, ?), and the amplitude of the X-ray variations is attributed to changes in the disk’s projected area and periodic obscuration by the outer disk (see Fig. 5A and Supplementary 6.1). The BH spin parameter can then be estimated by assuming that the observed period $T_{\rm prec}\sim 19.6$ days corresponds to the LT precession period. Adopting several power-law indices for the surface density profile (?), and a BH mass range of $M_{\bullet}=10^{6.7\pm 0.5}~M_{\odot}$, we found that the spin parameter $a_{\bullet}$ fell within the range of (0.460.14)-(0.46-0.14) or 0.110.350.11-0.35 (Fig. 6A). Since a negative spin parameter would result in a larger ISCO and, consequently, a larger inner disk radius, it would be challenging to explain the observed hot disk. Therefore, a positive spin parameter was favored. By modeling the X-ray lightcurve, we constrained the observer’s viewing angle relative to the BH spin to $\theta_{\rm obs}\sim 38.4^{+0.5\circ}_{-0.6}$, and the disk precession angle to $\theta_{\rm i}\sim 14.5^{\circ}\pm 0.5^{\circ}$ (Fig. 6B).
@@
We then examined the jet precession scenario, in which the radio luminosity varies as the jet cone shifts toward and away from the line of sight. The peak-to-dip ratio is determined by the Doppler factor ratio, which depends on the jet Lorentz factor ($\Gamma$) and the angle between the observer and the jet axis (Supplementary 6.2). The luminosity peaks when the difference between $\theta_{\rm obs}$ and $\theta_{\rm i}$ is minimized and dips when it is maximized. Adopting $\theta_{\rm obs}\sim 38.4^{\circ}$ and $\theta_{\rm i}\sim 14.5^{\circ}$, we derived $1.2\leq\Gamma\leq 1.6$. Jets can be powered by two primary mechanisms: the Blandford-Znajek (BZ; (?)) mechanism, driven by BH spin, and the Blandford-Payne (BP; (?)) mechanism, powered by disk rotation. Both mechanisms can account for our radio observations, with the BZ mechanism requiring a stronger magnetic field than the BP mechanism, i.e., $B_{\bullet}\sim 2.8\times 10^{3}\rm~G$ for the former and $B_{\bullet}\sim 1.5\times 10^{3}\rm~G$ for the latter, assuming a Lorentz factor of $\Gamma=1.6$. In Fig. 5B, we present a comparison between our disk-jet precession model and the observations.
@@
Discussion
Other potential mechanisms driving X-ray and radio co-variances
Disk radiation pressure instabilities may also produce comparable periodicity in X-ray and radio emissions. For example, the microquasar GRS 1915+105 exhibited simultaneous periodic modulations in X-ray and radio bands on timescales of 20–50 minutes (?, ?, ?). This periodic flaring activity has been attributed to radiation pressure instabilities, where material in the inner region of an optically thick accretion disk is rapidly depleted and replenished (?, ?). During this process, part of the inner disk is ejected to form a jet, as indicated by a rise in the radio flare coinciding with a dip in the X-ray flare (?). However, the X-ray and radio variations in GRS 1915+105 are considerably more complex, resulting in diverse correlations between the two-band lightcurves (?). Overall, the lack of simulations and theoretical models on the short-term evolution of jets during radiation-pressure instabilities makes it challenging to further test this scenario.
Nature of the 2024 re-brightening
[…]
In the latest NTT spectrum (2024-08-27, at +239 days), we detected high-ionization coronal lines (CLs). These include Fe xiv $\lambda 5303$, Fe vii $\lambda\lambda 5720,6087$, and Fe x $\lambda 6374$. There are indications of CLs in the first NTT spectrum taken before the seasonal gap, although they appear substantially weaker. CLs are high-ionization lines (e.g., of Iron, Neon, Argon, Sulfur) that originate from the photoionization or collisional ionization of a clumpy interstellar medium or other pre-existing material. Although typically associated with AGN, strong CLs have been detected in the optical spectra of some galaxies that show little to no evidence of AGN activity (?, ?). Many of these CLs have ionization potentials of $\geq 100$ eV, requiring a strong extreme UV (EUV) and/or soft X-ray ionizing continuum. It has long been thought that such extreme coronal line emitters (ECLEs) are powered by TDEs, and recent studies support this scenario (?, ?). Interestingly, a few optical-UV TDE candidates (?, ?) out of the $\sim 50$ strong candidates discovered to date have exhibited late ($\geq 200$ days post-peak) X-ray brightening accompanied by the emergence of iron coronal lines. In contrast, in the case of AT2020afhd, the CLs appear to emerge significantly later than the onset of X-ray emission. Further investigation, including detailed modeling, is required to understand the origin of the CLs, which will be explored in a forthcoming study.
@@ Discussion
Lomb-Scargle periodogram and cross-correlation function
The X-ray lightcurve was derived from the unabsorbed flux of the diskbb component in the 0.3–2 keV band, revealing clear quasi-periodic variations between August 3 and October 21, 2024, spanning a total of 79 days. To analyze these variations, we computed the LSP (?) using the astropy library (?) with data from this 79-day period. To more accurately determine the period of the variations, we divided the NICER observations into individual GTIs with durations exceeding 300 seconds. The flux for each NICER GTI was measured following the method described in Supplementary 1. Our analysis was performed on a GTI basis for the NICER data and on an observation basis for the XRT and PN data. Given the average sampling interval of 1 day and the total duration of 79 days, we conducted the LSP analysis over a period search range of 1 to 100 days, similar to the range used in the study of AT2020ocn, which had a period of 15 days (?). This analysis identified a period of 19.6 days with a FWHM of 3.0 days, as illustrated in the Fig. 3A.
To assess the statistical significance of the detected period, we first quantified the contribution of the LSP continuum by modeling it as a power-law, $P(f)\propto f^{\alpha}$. The fit was applied to the observed LSP, excluding the period range corresponding to the detected signal, as defined by its FWHM. The best-fitting power-law index was determined to be $\alpha=-0.02\pm 0.05$. This value is consistent with 0 within the $1\sigma$ uncertainty range, indicating that the continuum was not significantly affected by red noise but was instead dominated by white noise.
To further test whether the continuum was consistent with white noise, we applied the algorithms described in (?). We compared the cumulative distribution function (CDF) and probability density function (PDF) of the LSP power values to those expected for white noise. The expected CDF followed $1-\exp(-z)$ (?), where $z$ represented LSP powers. The comparisons of the observed and expected CDF and PDF are shown in Figs. 8A and B. To quantify these comparisons, we conducted Kolmogorov-Smirnov (K-S) and Anderson-Darling (A-D) goodness-of-fit tests on the observed and expected CDF using the scipy library (?). The K-S test yielded a statistic of 0.0391 (p-value = 0.93), and the A-D test returned a statistic of 0.4278, which is well below the 10% critical value of 1.07. These results indicate that, at a 90% confidence level, the null hypothesis that the LSP power followed the expected white noise distribution could not be rejected. Additionally, we performed 100,000 Monte Carlo simulations of the LSP continuum under the assumption of white noise and computed their corresponding K-S and A-D statistics. The distributions of these statistics are presented in Figs. 8C and D. The K-S and A-D statistic values of the observed CDF fell within the $1\sigma$ range of the distribution of the simulated white noise statistic values. These findings robustly indicated that the LSP continuum was statistically consistent with white noise.
To investigate the correlation between radio and X-ray variations of AT2020afhd, we used the Python-based discrete cross-correlation function (PyCCF, (?)), focusing on data collected between June 19 and October 21, 2024. The analysis revealed that the cross-correlation function (CCF) peaked at a time lag of $\sim-19.0$ days (see Fig. 3B). To assess the global significance of this correlation, we generated 100,000 random radio lightcurves, preserving the same observation sampling while randomizing the luminosities within the observed minimum and maximum values. We tested the CCF across time lags spanning 45-45 to $10$ days, a range covering $\pm 1$ cycle of the periodic signal. Among these simulations, only 2 spurious CCF signals were detected, corresponding to a statistical significance of approximately $4.26\sigma$. Furthermore, we estimated the uncertainties in the time lag by focusing on the range of 30-30 to 10-10 days, corresponding to a single period cycle around the CCF peak. Using 100,000 Monte Carlo simulations with a bin size of 0.1 days and bootstrapping methods to estimate uncertainties (?, ?), we determined a time lag of $19.0^{+0.7}_{-0.6}$ days for the radio relative to the X-ray. This result is consistent with the observed period, indicating that the X-ray and radio emissions were synchronized. As described in Section 1.3.5, we applied a 22% offset to the observed ATCA flux density. Using the original ATCA flux density and following the same methodology outlined above, we obtained a statistical significance of $3.85\sigma$ and a time lag of $19.1^{+0.6}_{-0.7}$ days. These results indicate that the calibration offset had no substantial impact on the derived time lag or the statistical significance of the correlation.